OpenStudy (anonymous):

hey guys, some one help me pls if a= b^n/c^m and the dimensions as follows a is [LT] b is [L^2*T^-1] c is [LT^2] find n and m . i reach to this solution [L^2*T^3]^m=[L^2*T^-1]^n

OpenStudy (anonymous):

L^2n+m T^-n+2m L^2n+m=L^1 T^-n+2m=T^1